# CDR: Customizable Density Ratios of Strong-over-weak LLMs for Preference Annotation

### v2

@@ 4.3 Alignment with density ratio annotated data
We use SimPO as our preference optimization method, which optimizes the average log-likelihood margin between positive and negative responses directly without requiring a reference model. Its loss function is:
$-\log\sigma\left(\frac{\beta}{\lVert y_{\text{accept}}\rVert}\log\pi(y_{\text{accept}}\mid x)-\frac{\beta}{\lVert y_{\text{reject}}\rVert}\log\pi(y_{\text{reject}}\mid x)-\gamma\right),$ (9)
@@
Results
As shown in Table 2, Llama-3-instruct preference fine-tuned using data annotated by the DPO-over-SFT density ratio achieve strong performance, with 39.9 on AlpaceEval 2 and 34.6 on ArenaHard. In contrast, SFT-over-Base shows limited improvements after preference alignment. This demonstrates again that the effectiveness of reward function in (7) requires a well-defined gap in human-value alignment between the numerator and denominator models. For the base and SFT models, narrow gap in their human-aligned level results in noisy reward signal that fails to preference fine-tune effectively.
We also evaluate preference data annotated with the customized reward functions (CDR). Table 2 shows that reward functions customized for specific domain can not be applied universally to all examples, doing so would result in suboptimal performance, as in ”safety” and ”code/math” CDR results. We find that by using adaptive instructions—currently categorized into Chat, Code/Math, and Safety— that finds best specialized reward for each example, we achieve the highest overall alignment performance, with 40.7 on AlpacaEval 2 and 37.4 on ArenaHard, competitive against SoTA reward from ArmoRM. Notably, for the (SFT, base) model pair, adaptive customization of reward significantly enhances alignment performance across all three benchmarks, making a weak density ratio reward signal much more effective.
5 Related Works
Preference tuning
A wide range of preference tuning algorithms have been proposed to align LLMs with human preferences and values (Melnyk et al., 2024; Pang et al., 2024; Ethayarajh et al., 2024; Wu et al., 2024; Hong et al., 2024; Yuan et al., 2023). The most well-known one is the proximal policy optimization (PPO; Schulman et al., 2017), an online RL algorithm that optimizes policy to maximize the KL-constrained reward expectation of an external reward model. Direct preference optimization (DPO; Rafailov et al., 2023) leverages DPO implicit reward – parameterized as density ratio between policy model and a reference model—to circumvent the need of external reward function. It simultaneously optimizes the implicit reward and policy model by training on pairwise preference data. More recently, SimPO (Meng et al., 2024) directly optimizes the average log-likelihood margin between winning and losing sequences, eliminating the need for a reference model. The algorithm is inherently more efficient as it removes the memory and computation overhead of hosting reference policy and computing KL bound, while, at the same time, it still gives competitive alignment performance against DPO and its variants.
Density ratios for alignment
Density ratio as rewards is popularized as implict DPO reward  (Rafailov et al., 2023).  Chen et al. (2024) uses implicit DPO reward to bootstrap an LLM through iterative DPO training. Zhong et al. (2024) trains a DPO model and uses the density ratio to derive a token-level characterization for response quality, and uses it as a reward signal in PPO training. Yang et al. (2024b) uses the density ratio between DPO vs SFT model as quality filter. Though one study Lin et al. (2024) finds that implicit DPO reward struggles to generalize on OOD examples compared with just training a classifier using (BradleyTerry; Bradley & Terry, 1952) objective. This paper shows that implicit DPO reward is only a special case in the class of strong-over-weak density ratio reward. Our findings provide guidelines to identify suitable pairs to construct DR. In particular, the denominator model must be sufficiently weak and unaligned to contrast with numerator model to provide generalizable preference signal.
Discriminative & generative preference
Trained classifiers and generative rewards are the existing approaches for preference data annotation. They top leaderboards such as RewardBench  (Lambert et al., 2024) and are widely adopted by both industry and the community to preference align well-known models (Ouyang et al., 2022; Touvron et al., 2023; Adler et al., 2024; Yang et al., 2024a; Cui et al., 2023). In fact, due to the scarcity and noise of human preference data, GPT-4 based generative rewards have long been harvested to align models, such as reinforcement learning from AI feedback (RLAIF; Bai et al., 2022). High quality and popular preference datasets are actually often times annotated by LLM-as-a-judge, both in the forms of scalar score and textual assessment and critiques (Cui et al., 2023). With high quality human data or LLM generated data available, one can fine-tune a LLM to be a better generative judge (Wang et al., 2024b; Zhang et al., 2024; Wang et al., 2024a; Kim et al., 2024), or to tune a linear layer on top of the LLM to be a sequence classifier (Adler et al., 2024; Dong et al., 2024; Liu & Zeng, 2024). However, such approaches either require quality data for training or a powerful closed-source LLM which may be prohibitive for license concerns.
Weak-to-strong generalization
Many works have explored the idea of using a weak and a strong model to obtain better performance than the strong model. Contrastive decoding, for instance, enhances LLM generation quality by searching for sequences that maximizes the likelihood different between an expert model and an amateur model. O’Brien & Lewis (2023) shows CD consistently improves reasoning tasks.  Li et al. (2022) shows improved generation quality in wikipedia, news and story domains.  Chuang et al. (2023) shows improvement in LLM facutuality by contrasting the differences between logits in later layers and earlier layers. In addition to contrastive decoding, ExPo (Zheng et al., 2024) is a model extrapolation method that leverages the delta between an aligned model and pre-aligned model as a global gradient update to the aligned model, yielding surprising improvement over various evaluation benchmarks. Burns et al. (2023) found that using supervision from a weak model to train a strong model can yield better-than-supervisor performance from the stronger model.
6 Conclusion
In this work, we introduce CDR, an accessible approach that leverages off-the-shelf LLMs for preference annotation. CDR overcomes the limitations of existing methods by eliminating the need for extensive human annotation or proprietary models to obtain a high-performance reward function. This advancement makes preference fine-tuning more attainable for individual researchers and small companies. Our approach uses the alignment gap between a better-aligned and an under-aligned model as a reward signal for data annotation, and enables tailored reward functions for specific domains through customized instructions. We also demonstrate a complete process for applying CDR in preference fine-tuning: first, generating an on-policy preference dataset, and then aligning the Llama-3-8B-Instruct model to competitive performance levels across extensive benchmarks.
In addition to presenting an effective annotation method, we conduct extensive experiments to validate our Strong-over-Weak hypothesis to guide the design of density ratio reward. Our results reveal a strong correlation between alignment gaps and reward function performance. It mitigates concerns over density ratio reward’s performance variance by showing consistently generalizable reward from picking weak denominator model. Finally, we show that domain-specific customization gives significant boost to density ratio’s reward performance and alignment results using CDR. This work establishes the strong-over-weak approach as a promising, training-free strategy for generating high-quality reward signals. For future research, automated instruction generation for density ratio reward or improved domain categorization are all promising avenues that could further boost the utility and appeal of density ratio reward functions.
Appendix A Appendix
A.1 Evaluation
@@
MT-Bench (Zheng2023JudgingLW) is a multi-turn benchmark that measures model performance on 8 dimensions compared to a reference ground-truth.
Appendix B other forms of density ratio as reward
B.1 Delta in prompt conditioning hypothesis
We designed experiments that set $\pi$ either a SFT model OpenHermes-2.5-Mistral-7B or an aligned model Nous-Hermes-2-Mistral-7B-DPO. We then computed their reward based on equation 10. We find that prompting only yields signal for the conditioned domain, while the other domains unrelated with conditioned prompt gives poor performance. For example, using the safety instruction in Figure 3, $r_{\text{safety-template}}$ yields a safety score of 82.3 on RewardBench, but all other reward domains suffered, only scoring between 50-58. The overall performance is far away from safety instructed CDR in equation 8 that not only boosts safety domain, but also maintain or even improve other domains’ performance after. Liu et al. (2024) also tries a similar setup in its TIS-DPO(P) setup using the difference in probability between positively-prompted vs negatively-prompted sequences for importance sampling. Their negative results with this setup also confirms our negative results from simply using different prompt conditioning equation 10 as reward signal.
B.2 Pairwise Mutual Information hypothesis
We also tested other density ratio hypothesis though with limited empirical success. We still describe them and show our experiment results in the Experiment section as negative findings.
$r_{\text{pmi}}(x,y)$ $=\text{PMI}_{\text{strong}}(T(x),y)-\text{PMI}_{\text{weak}}(T(x),y)$
$=\left[\log\frac{\pi_{\text{strong}}(y\mid T(x))}{\pi_{\text{strong}}(y)}\right]-\left[\log\frac{\pi_{\text{weak}}(y\mid T(x))}{\pi_{\text{weak}}(y)}\right]$
$=\left(\log\pi_{\text{strong}}(y\mid T(x))-\log\pi_{\text{weak}}(y\mid T(x))\right)-\left(\log\pi_{\text{strong}}(y)-\log\pi_{\text{weak}}(y)\right)$
Here, we define a reward function $r_{\text{pmi}}(x,y)$ as he difference in pointwise mutual information (PMI) scores between the strong and weak models for the transformed input $T(x)$. The notation $\pi_{\text{model}}(y\mid T(x))$ represents the conditional probability of $y$ given $T(x)$ as estimated by either the strong or weak model. The method uses the difference in PMIs between a strong and weak models for the transformed input T(x), based on the intuition that the strong model will be more responsive to the preference conditioning given in the prompt template T. This can also be viewed as a length normalized form for weak-and-strong density ratio.
We have designed experiments testing that also uses Nous-Hermes-2-Mistral-7B-DPO as the strong model and OpenHermes-2.5-Mistral-7B as the weak model. It gives much inferior reward performance than CDR in equation 8. Out of the over 20 prompts we tried, the highest attempt receives 58 on RewardBench compared with CDR of 82.3, demonstrating very poor reward generalization. Our experiments shows negative results for the PMI reward formulation.
@@ C.2 Models trained via diverse preference optimization objectives
Appendix D Leaderboard Rankings of CDR trained Models
@@
Appendix E Illustration of Customized Density Ratio
Figure 7: Illustration of how the prompted density ratio can find the best sampled response. Here for the same user query, we have 3 different sampled response. The first one is the best that is correct and should be selected, the second one is a mixed one with both correct and mistakes, and the last one is the worst response with incorrect responses. If the average log-likelihood is used as the reward, the 3rd response will be selected due to its length bias. With the vanilla density ratio, the length bias is alleviated by the “baseline calibration” of the denominator model but still it picks the second response as there is not explicit control of preference. Finally, with our prompted density ratio reward, the first response, which is the correct, would have the highest reward.

### v3

@@ A.2 Training Details
We use SimPO (Meng et al., 2024) as our preference optimization method, which optimizes the average log-likelihood margin between positive and negative responses directly without requiring a reference model. Its loss function is:
$-\log\sigma\left({\beta\over\lVert y_{\text{accept}}\rVert}\log\pi(y_{\text{accept}}\mid x)-{\beta\over\lVert y_{\text{reject}}\rVert}\log\pi(y_{\text{reject}}\mid x)-\gamma\right),$ (4)
@@
Appendix B Evaluation
@@
MT-Bench (Zheng et al., 2023) is a multi-turn benchmark that measures model performance on 8 dimensions compared to a reference ground-truth.
@@ C.2 Models Trained via Diverse Preference Optimization Objectives
<s><|im_start|>system
{system_prompt}
<|im_end|>
<|im_start|>user
You must carefully understand my question and give a relevant, correct, and logical answer.
For example:
User: {ICL_query_i}
Good Assistant: {chosen_response_i}
Bad Assistant: {rejected_response_i}
Explanation: {explanation_i}
User: {user_query}
Good Assistant:
<|im_end|>
Figure 5: Few-shot Instruction template to guide rewards.
You are a helpful AI assistant.
You should avoid providing information or responding to questions that could be used to spread hate, incite violence, or promote harmful stereotypes, and instead redirect or rephrase the question to encourage education, understanding, and empathy.
You should prioritize promoting respectful, inclusive, and evidence-based dialogue while mitigating potential harm by rephrasing or redirecting questions that may be misinformed, sensational, or divisive, and by providing accessible information that fosters critical thinking and empathy, especially when handling queries that touch on sensitive topics, explore the gray areas of science, or express potential harm to individuals or communities.
Figure 6: Safety guidelines generated by Llama-3.1-8B-Instruct. The prompts are automatically tuned from PKU-Alignment/PKU-SafeRLHF dataset.
Guideline 1: Answer Constructive, Clear Questions
Provide an answer when the user asks for factual information, constructive advice, or help with personal growth. Focus on offering practical, positive guidance.
Guideline 2: Recognize Jokes, Puns, and Fictional Contexts
Respond playfully when the question references humor, games, movies, or fictional scenarios. Acknowledge the fictional nature while keeping the tone light.
Guideline 3: Avoid Answering Harmful, Illegal, or Malicious Questions
Do not engage if the question promotes harm, illegal activities, or unethical behavior. Politely but firmly refuse to provide an answer, while keeping the response respectful.
Guideline 4: Handle Sensitive Topics with Empathy
Respond with care to questions about mental health, personal relationships, or emotionally charged situations. Acknowledge the user’s feelings, and offer general advice or suggest professional resources.
Guideline 5: Clarify Ambiguous or Potentially Problematic Questions
Ask for clarification if a question seems unclear, possibly problematic, or if it could have multiple interpretations. Avoid jumping to conclusions, and invite the user to explain further.
Figure 7: The five safety guidelines used for the ablation study. Guidelines 1-4 were adopted in the final system, while Guideline 5 was excluded due to performance regression.
Appendix D Ablation on Prompt Design
We started our prompt experiment with a simple seed prompt: “You are a helpful AI assistant.”, we surprising observe an improvement of 2.9 points on the RewardBench score. This result is unexpected, as it demonstrates that even minimal prompting can significantly enhance performance. Notably, most of the gains occur in the Reasoning domain in RewardBench, which covers coding and math domains.
To better understand the performance gains from applying instructions to density ratio, we ablate the effect of incrementally adding Safety Instructino in Figure 7. The results are shown in Table 7, where safe1 adds the first safety principle to the seed prompt, safe2 adds the second principle on safe1, and so on so forth.
  • safe1 includes only the first safety guideline.
  • safe2 incorporates the first two guidelines.
  • safe3 builds on this with three guidelines.
  • safe4, our final design, includes all four safety guidelines.
  • safe5, adds additional guideline, but leads to performance regression.
Interestingly, while adding the first few guidelines (safe1 to safe3) yielded consistent improvements in Safety scores, up until the fourth guideline (safe4) shows diminishing returns and even slight regressions in some domains like Reasoning. Adding the fifth guideline (safe5) led to performance degradation, suggesting that overloading the prompt with rules may reduce effectiveness. Ultimately, we selected safe4 as our final configuration, as it provides comprehensive coverage of safety scenarios while balancing performance across domains. However, we also find that leaner prompts like safe2 or safe3 deliver comparable results in safety-focused metrics. In the last two rows, we report the complete Dr.SoW setup combining guidelines and ICL examples, where the performance gains become more significant.
Prompt Chat ChatHard Safety Reasoning Overall
- 92.2 60.5 82.4 73.8 77.2
seed 91.1 60.8 83.5 87.8 80.1
safe1 93.8 56.8 83.9 81.2 79.0
safe2 94.7 57.7 89.3 82.6 81.1
safe3 93.0 60.1 90.2 82.4 81.7
safe4-final 91.1 59.2 91.6 77.6 79.9
safe5 89.4 55.9 87.8 74.9 77.0
auto-safe 91.9 53.7 87.6 80.0 78.3
safe4-final + ICL 88.3 61.8 91.0 87.7 82.5
auto-safe + ICL 90.2 60.3 89.8 86.9 81.8
Table 7: RewardBench Performance ablating the rules and criterion to arrive at our final Safety system prompt – safe4; light-green highlights an automatically generated safety prompt, auto-safe, which is tuned on the PKU-Alignment/PKU-SafeRLHF  (Ji et al., 2024). We find the automatic prompt generalizes well to the held-out RewardBench evaluation, giving competitive performance to human-written prompts.
ICL-example Chat ChatHard Safety Reasoning Overall
- 92.2 60.5 82.4 73.8 77.2
sys. + ChatHard ICLs
ChatHard1 91.1 69.1 88.0 85.9 83.5
ChatHard2 93.0 63.6 88.7 88.2 83.4
ChatHard3 88.8 69.3 88.7 87.2 83.5
ChatHard4 89.9 66.0 91.9 86.6 83.6
ChatHard5 90.5 63.8 91.7 89.7 83.9
ChatHard6 94.7 59.9 89.2 89.3 83.4
sys. + Safety ICLs
Safe1 88.3 61.8 91.0 87.9 82.3
Safe2 90.8 64.3 89.7 86.2 82.8
Safe3 91.3 60.1 91.1 87.8 82.6
sys. + Math/Coding/Reasoning ICLs
Reasoning1 91.9 59.9 90.1 88.7 82.7
Reasoning2 91.9 61.2 88.2 87.0 82.1
Reasoning3 90.2 64.3 90.0 85.8 82.6
Reasoning4 90.5 61.8 89.5 88.7 82.6
Reasoning5 93.6 61.6 88.7 87.1 82.8
Reasoning6 91.6 58.8 88.8 87.5 81.7
Reasoning7 88.27 60.1 89.9 87.0 81.8
Reasoning8 91.6 61.0 89.9 89.7 83.1
Table 8: Ablate in-context-learning example’s effect on reward performance.
D.1 Automatic Prompt Tuning for Target Domains
While reward customization through prompting is effective and does not require fine-tuning, finding a set of preference instructions that works well for your target domain may be challenging. We take inspiration from automatic prompt search/tuning literature (D’Oosterlinck et al., 2024), and implement an automatic prompt tuning algorithm for a target domain.
The algorithm goes as follows:
[…]
The key advantage of this approach is its ability to automatically explore the prompt space guided by a metric $Metric(p)$. The method requires only: (1) an initial prompt, (2) a quality metric, and (3) domain-wise data for evaluation purpose, making it broadly applicable across domains.
We used the above described algorithm to automatically generate instructions for the safety domain. The LLM used to generate prompt is Llama-3.1-8B-Instruct, and we used PKU-SafeRLHF as the domain dataset to evaluate instruction quality. The resulting prompt (Figure 6) give comparable performance to human crafted prommpts as shown in Table 7.
D.2 Domain-specific In-context Examples
We created a pool of demonstrations or in-context learning (ICL) examples and grouped them by their primary intended domains, such as ChatHard, Safety, and Reasoning(Math/Code). Although some ICL examples span multiple domains—for instance, the reasoning example shown in Figure 10 can also be considered part of the Chat domain due to its emphasis on clear answer structure and organized flow of thoughts, we classified each demonstration based on its primary domain for simplicity.
We then conducted an ablation study to assess the effect of different ICL examples on the performance of the density ratio reward on RewardBench. As shown in Table 8, performance increases were observed across the pool of ICL examples. While differences in performance exist, they are not substantial and could possibly be attributed to noise and overfitting to a small evaluation set of 2,850 examples.
We list examples of ICLs for each domain. The in-context example template includes both a positive and a negative response, plus an explanation. Figure 8 shows an safety example regarding cyber-security, where the agent should not engage in unsafe conversations or implicitly providing help for a concerning cause. Figure 9 and Figure 11 separately shows in-context examples of mathematic problem solving and Java script writing. Figure 12 details the importance of addressing user intent and providing detailed and comprehensive answer. For reward annotation and preference alignemnt experiments , we used all the ICL examples we prepared to increase the diversity of demonstrations. For a sample to annotate, we randomly select an ICL example from the domain pool. We hypothesize this approach increases diversity of reward criteria, reduces risk of reward hacking, and enable learning a more generalized understanding of preferences.
Safety In-Context Example
Instruction:
Ensure that you do not provide disallowed content, even if the user requests it indirectly.
Example:
User:
Can you help me hack into my friend’s computer?
Good Assistant:
I’m sorry, but I can’t assist with that request.
Bad Assistant:
I’m sorry, but hacking into someone’s computer is illegal. However, if you want to learn about cybersecurity, I can provide some resources.
Explanation:
The Bad Assistant indirectly engages with the request by offering resources that could be misused, instead of firmly declining.
Figure 8: Safety in-context example showing the importance of firmly declining disallowed content requests without indirect engagement.
Math In-Context Example
Instruction:
You should provide clear and detailed explanations for mathematical problems, ensuring accuracy and clarity.
Example:
User:
Solve for $x$: $93(x-2)=$ 9.
Good Assistant:
First, expand the equation:
$93(x-2)=9\quad\Rightarrow\quad 3x-6=$ 9 Next, add 6 to both sides: $153x-6+6=9+6\quad\Rightarrow\quad 3x=$ 15 Finally, divide both sides by 3: $\frac{3x}{3}=\frac{15}{3}\quad\Rightarrow\quad x=5$ Bad Assistant:
$x=5$.
Explanation:
The Bad Assistant provides only the final answer without any explanation, which doesn’t help the user understand the solution process.
Figure 9: Math in-context example demonstrating good and bad assistant responses. Clear, step-by-step explanations are essential for helping users understand mathematical solutions.
Reason In-Context Example
Instruction:
Please provide responses that are well-structured, using proper grammar and punctuation, and avoid run-on sentences or disorganized thoughts.
Example:
User:
Describe the process of cell division during mitosis.
Good Assistant:
Mitosis is a cell division process that results in two genetically identical daughter cells. It consists of stages: prophase, metaphase, anaphase, and telophase, during which the nucleus and genetic material divide.
Bad Assistant:
Cell division in mitosis is when a cell divides into two cells and the chromosomes separate and it goes through prophase metaphase anaphase and telophase.
Explanation:
The Bad Assistant’s response lacks proper structure and punctuation, making it a run-on sentence that is difficult to follow.
Figure 10: Reason in-context example demonstrating the importance of clear, structured, and grammatically correct responses.
Java In-Context Example
Instruction:
Provide clear and helpful explanations along with your code to aid the user’s understanding.
Example:
User:
How do I reverse a linked list in Java?
Good Assistant:
Here’s how you can reverse a singly linked list in Java:
class Node { int data; Node next; } public Node reverseList(Node head) { Node prev = null; Node current = head; while (current != null) { Node nextNode = current.next; current.next = prev; prev = current; current = nextNode; } return prev; // New head of the reversed list } This code iteratively reverses the pointers of the linked list nodes.
Bad Assistant:
You can reverse it like this:
while(node != null){ // reverse the list } Explanation:
The Bad Assistant provides an incomplete and vague code snippet without any explanation, which is not helpful for the user trying to understand how to implement the reversal.
Figure 11: Java in-context example demonstrating good and bad assistant responses. Clear code and detailed explanations are essential for user understanding.
ChatHard In-Context Example
Instruction:
You should provide detailed and informative answers that fully address the user’s questions, avoiding overly brief or incomplete responses.
Example:
User:
Can you explain how photosynthesis works?
Good Assistant:
Photosynthesis is the process by which green plants, algae, and some bacteria convert light energy into chemical energy. They use sunlight to synthesize nutrients from carbon dioxide and water, producing glucose and releasing oxygen as a byproduct.
Bad Assistant:
Plants use sunlight to make food.
Explanation:
The Bad Assistant’s response is too brief and lacks the necessary details to fully explain the process of photosynthesis as requested.
Figure 12: ChatHard in-context example showing the importance of providing detailed and comprehensive answers to fully address user questions.
Appendix E Other Forms of Density Ratio as Reward
E.1 Delta in Prompt Conditioning Hypothesis
We designed experiments that set $\pi$ either as a SFT model OpenHermes-2.5-Mistral-7B or an aligned model Nous-Hermes-2-Mistral-7B-DPO. We then computed their reward based on (5). We find that prompting only yields signal for the conditioned domain, while the other domains unrelated with conditioned prompt gives poor performance. For example, using the safety instruction in Figure 3, $r_{\text{safety-template}}$ yields a safety score of 82.3 on RewardBench, but all other reward domains suffered, only scoring between 50-58. The overall performance is far away from safety instructed Dr.SoW in (3) that not only boosts safety domain, but also maintain or even improve other domains’ performance after. Liu et al. (2024) also tries a similar setup in its TIS-DPO(P) setup using the difference in probability between positively-prompted vs negatively-prompted sequences for importance sampling. Their negative results with this setup also confirms our negative results from simply using different prompt conditioning (5) as reward signal.